Solve:

Solve:

u_t-u_x_x=e^{-t}, x\epsilon R, t>0 \\ u(x,0)=e^{-x}, x\epsilon R
Answer #1

rae wh Du ut at Civen PDE- ut - 4 xx = tet (0 - Di?) u = text - where ay - Du Mere D-Ds is not a linear Uxx 2 = D²u factor iParticular integral Pisaés D-D 2 Elttoa sina att be IF (DD) at & be (-1) frantong given blince general solepsion of differentL.I. so that Theorem. To show that F(D, D) eax+by = F(a,b) ear+by Proof. We know that F(D, D) consists of terms of the form